### v2

# Grothendieck inequalities characterize converses to the polynomial method

Abstract.
A surprising ‘converse to the polynomial method’ of Aaronson et al. (CCC’16) shows that any bounded quadratic polynomial can be computed exactly in expectation by a 1-query algorithm up to a universal multiplicative factor related to the famous Grothendieck constant. Here we show that such a result does not generalize to quartic polynomials and $2$-query algorithms, even when we allow for additive approximations. We also show that the additive approximation implied by their result is tight for bounded bilinear forms, which gives a new characterization of the Grothendieck constant in terms of $1$-query quantum algorithms. Along the way we provide reformulations of the completely bounded norm of a form, and its dual norm.
This version extends arXiv version 1 of this work with most of [BEG22], which appeared in the proceedings of TQC’22.
[Uncaptioned image] This research was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 945045, and by the NWO Gravitation project NETWORKS under grant no. 024.002.003.
1. Introduction
Quantum query complexity is one of the few models of computation in which the strengths and weaknesses of quantum computers can be rigorously studied with currently-available techniques (see e.g., [Amb18, Aar21] for recent surveys). On the one hand, many of quantum computing’s best-known algorithms, such as for unstructured search [Gro96], period finding (the core of Shor’s algorithm for integer factoring) [Sho97] and element distinctness [Amb07], are most naturally described in the query model. On the other hand, the model admits powerful lower-bound techniques.
For a (possibly partial) Boolean function $f:D\to\{-1,1\}$ defined on a set $D\subseteq\{-1,1\}^{n}$, the celebrated polynomial method of Beals, Buhrman, Cleve, Mosca and de Wolf [${}^{+}$01] gives a lower bound on the quantum query complexity of $f$, denoted $Q(f)$, in terms of the minimal degree of an approximating polynomial for $f$, or approximate degree, $\widetilde{\deg}(f)$. The method relies on the basic fact that for any $t$-query quantum algorithm $\mathcal{A}$ that takes an $n$-bit input and returns a sign, there is a real $n$-variable polynomial $p$ of degree at most $2t$ such that $p(x)=\mathbb{E}[\mathcal{A}(x)]$ for every $x$. Here, the expectation is taken with respect to the randomness in the measurement done by $\mathcal{A}$ [^1] We identify quantum query algorithms with the (random) functions giving their outputs. Using this method, many well-known quantum algorithms were proved to be optimal in terms of query complexity (see e.g., [BKT20] and references therein).
Since polynomials are simpler objects than quantum query algorithms, it is of interest to know how well approximate degree approximates quantum query complexity. There are total functions $f$ that satisfy $Q(f)\geq\widetilde{\deg}(f)^{c}$ for some absolute constant $c>1$ [Amb06, ABDK16]; the second reference gives an exponent $c=4-o(1)$, which was shown to be optimal in [ABDK16]. For partial functions it was recently shown that this separation can even be exponential [AB23]. These separations rule out a direct converse to the polynomial method, whereby a given bounded degree-$2t$ polynomial $p$ can be computed by a $t$-query quantum algorithm $\mathcal{A}$. However, since these results concern functions whose approximate degree grows with $n$, they leave room for the possibility that such an $\mathcal{A}$ approximates $p$ with some error that depends on $t$.
We will say that a polynomial $p$ is bounded if its restriction to the Boolean hypercube takes values in the interval $[-1,1]$ and that $\mathcal{A}$ approximates $p$ if for some constant additive error parameter $\varepsilon<1$, we have that $|p(x)-{\mathbb{E}}[\mathcal{A}(x)]|\leq\varepsilon$ for every $x$. Note that an additive error of 1 can trivially be achieved with a uniformly random coin flip. For a function $p:\{-1,1\}^{n}\to\mathbb{R}$ and positive integer $t$, we denote the smallest additive error that a $t$-query quantum algorithm can achieve by
(1) $\mathcal{E}(p,t):=\inf\big{\{}\varepsilon\geq 0\,\mid\,$ $\ \exists\ t\text{-query quantum algorithm }\mathcal{A}\text{ with }$
$\ |p(x)-{\mathbb{E}}[\mathcal{A}(x)]|\leq\varepsilon\quad\forall x\in\{-1,1\}^{n}\big{\}}.$
For bounded polynomials of degree at most 2, a “multiplicative converse” to the polynomial method was proved in  [${}^{+}$16], showing that up to an absolute constant scaling, quadratic polynomials can indeed be computed by 1-query quantum algorithms.
Theorem 1.1 (Quadratic multiplicative converse [${}^{+}$16]).
There exists an absolute constant $C\in(0,1]$ such that $\mathcal{E}(Cp,1)=0$ for every bounded polynomial $p$ of degree at most $2$.
This result directly implies the following additive version.
Corollary 1.2 (Quadratic additive converse).
There exists an absolute constant $\varepsilon\in(0,1)$ such that the following holds. For every bounded polynomial $p$ of degree at most $2$, we have $\mathcal{E}(p,1)\leq\varepsilon$. In particular, one can take $\varepsilon=1-C$ for the constant $C$ appearing in Theorem 1.1.
In light of the polynomial method, Corollary 1.2 shows that one-query quantum algorithms are roughly equivalent to bounded quadratic polynomials. The authors of [${}^{+}$16] asked whether this result generalizes to higher degrees. Two ways to interpret this question are that for any $k$, any bounded degree-$2k$ polynomial $p$ satisfies:
  1. (a)
    Multiplicative converse: $\mathcal{E}(Cp,k)=0$ for some $C=C(k)>0$, or;
  2. (b)
    Additive converse: $\mathcal{E}(p,k)\leq\varepsilon$ for some $\varepsilon=\varepsilon(k)<1$.
The dependence on the degree $k$ in these options is necessary due to the known separations between bounded-error quantum query complexity and approximate degree. Option (a), the higher-degree version of Theorem 1.1, was ruled out in [ABP19].
Theorem 1.3 (Arunachalam–Briët–Palazuelos).
For any $C>0$, there exist an $n\in\mathbb{N}$ and a bounded quartic $n$-variable polynomial $p$ such that no two-query quantum algorithm $\mathcal{A}$ satisfies ${\mathbb{E}}[\mathcal{A}(x)]=Cp(x)$ for every $x\in\{-1,1\}^{n}$.
The original proof of this result was based on a probabilistic example, but that proof turns out to contain an error. Here, we show that Theorem 1.3 holds as stated, both by considering a slightly modified probabilistic example and by giving a completely explicit example. This result thus shows that the hoped-for constant $C(4)$ does not exist. This does not exclude the possibility of option (b), namely that Corollary 1.2 generalizes to higher degrees.
Our main result shows that it does not.
Theorem 1.4.
There is no constant $\varepsilon\in(0,1)$ such that for every bounded polynomial $p$ of degree at most $4$, we have $\mathcal{E}(p,2)\leq\varepsilon.$
In the context of quantum query complexity of Boolean functions, this rules out arguably the most natural way to upper bound $Q(f)$ in terms of $\widetilde{\deg}(f)$: First, $\varepsilon$-approximate $f$ by a degree-$2t$ polynomial $p$, then $\varepsilon^{\prime}$-approximate $p$ with a $t$-query quantum algorithm $\mathcal{A}$, with $\varepsilon+\varepsilon^{\prime}<1$, and finally boost the success probability of $\mathcal{A}$ so that it approximates $f$, for instance by taking the majority of independent runs of $\mathcal{A}$. Corollary 1.2 gives the only exceptional case where this is possible in general.
1.1. Best constants
Theorem 1.1 was proved using a surprising application of the famous Grothendieck theorem from Banach space theory [Gro53] (see also Section 1.2). For bounded bilinear forms, $p(x,y)=x^{\mathsf{T}}Ay$ given by a matrix $A\in\mathbb{R}^{n\times n}$, the result holds with $1/C$ equal to the Grothendieck constant $K_{G}$ (see [ABP19, Section 5] for a short proof). Determining the precise value of $K_{G}$ is a notorious open problem posed in [Gro53]; the best-known lower and upper bounds place it in the interval $(1.676,1.782)$ [Dav84, Ree91, BMMN13]. The general form of Theorem 1.1 follows from decoupling techniques. It is not hard to show that $1/K_{G}$ is the optimal constant in the bilinear case for the multiplicative setting of Theorem 1.1. Here, we also show that the additive approximation implied by the multiplicative setting is optimal.
Theorem 1.5.
The worst-case minimum error for one-query quantum algorithms satisfies
$\sup_{p}\mathcal{E}(p,1)=1-\frac{1}{K_{G}},$
where the supremum is taken over the set of bounded bilinear forms.
Theorem 1.5 may open a new route to determine the value of $K_{G}$ by studying the power and limitations of 1-query quantum algorithms. This complements another well-known characterization of $K_{G}$ in terms of the largest-possible Bell-inequality violations in two-player XOR games [Tsi80].
1.2. Techniques
The starting point for this work is a characterization of quantum query algorithms with completely bounded polynomials [ABP19]. Recently, this characterization was also used to make progress on the problem to determine “the need for structure in quantum speed-ups” in works by Bansal, Sinha and de Wolf [BSdW22] and by the second author [Esc23]. In addition, it led to a new exact SDP-based formulation for quantum query complexity, due to Laurent and the third author [GL19].
Here, we use this characterization to show that for a special class of polynomials of degree $2t$, the parameter $\mathcal{E}(p,t)$ can be written in terms of a ratio of norms related to norms appearing naturally in Grothendieck’s theorem. This special class of polynomials consists of block-multilinear forms, which have appeared naturally in the polynomial method and play an important role in other works related to quantum query complexity [OZ15, ${}^{+}$16, BSdW22], theoretical computer science [KN07, Lov10, KM13] and in the polarization theory of functional analysis [BH31, Har72]. To illustrate our approach, in the remainder of this introduction we focus on the quadratic case and the connection of our work to Grothendieck’s theorem; the quartic case is briefly touched on in the end of this introduction.
For the purpose of this section, we focus for a moment on the class of $2n$-variable bilinear polynomials of the form $p(x,y)=x^{\mathsf{T}}Ay$ for a matrix $A\in\mathbb{R}^{n\times n}$ A central role is played by two norms associated to $A$ (and hence $p$). The first is the $\ell_{\infty}$-norm of $A$ when identified with the map $(x,y)\mapsto p(x,y)$,
$\|A\|_{\infty}=\max_{x,y\in\{-1,1\}^{n}}x^{\mathsf{T}}Ay.$
The second is its completely bounded norm, which may be given by
(2) $\|A\|_{{\mathrm{cb}}}=\sup_{\begin{subarray}{c}d\in\mathbb{N},u,v:[n]\to S^{d-1}\end{subarray}}\sum_{i,j=1}^{n}A_{i,j}\langle u(i),v(j)\rangle,$
where $S^{d-1}$ denotes the real $(d-1)$-dimensional Euclidean unit sphere. The celebrated Grothendieck theorem asserts that these norms are equivalent up to a constant factor.
Theorem 1.6 (Grothendieck’s theorem).
There exists a constant $K<\infty$ such that for any $n\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$, we have
(3) $\|A\|_{\infty}\leq\|A\|_{{\mathrm{cb}}}\leq K\|A\|_{\infty}.$
The non-trivial part of this result is the second inequality in (3) and the smallest $K$ for which it holds is the above-mentioned Grothendieck constant $K_{G}$. Our characterization of $\mathcal{E}(p,1)$ involves the dual norms of the $\ell_{\infty}$-bilinear norm and the completely bounded norm (c.f. Section 2.1). The dual formulation of (the second inequality in) Grothendieck’s theorem asserts that, for any matrix $A$,
$\|A\|_{\infty,*}\leq K_{G}\|A\|_{{\mathrm{cb}},*}.$
Our characterization of $\mathcal{E}(p,1)$ can be stated as follows. Here $\langle A,B\rangle$ denotes the usual trace inner product on matrices.
Theorem 1.7 (Bilinear version of Theorem 5.1).
For a bounded bilinear polynomial $p$ of the form $p(x,y)=x^{\mathsf{T}}Ay$ for some matrix $A\in\mathbb{R}^{n\times n}$, we have
$\mathcal{E}(p,1)=\sup_{B\in\mathbb{R}^{n\times n}}\frac{\langle A,B\rangle-\|B\|_{{\mathrm{cb}},*}}{\|B\|_{\infty,*}}.$
In Section 7 we show how Theorem 1.5, our characterization of $K_{G}$ in terms of quantum query complexity, follows from this by maximizing over $A$.
We prove Theorem 1.4 by combining our main technical result, Theorem 5.1, which gives an analogous statement to Theorem 1.7 for higher degree polynomials, and one of the constructions behind Theorem 1.3.
2. Preliminaries
2.1. Notation
For $n\in\mathbb{N}$, write $[n]:=\{1,\dots,n\}$. We endow $\mathbb{R}^{d}$ with the standard inner product $\langle x,y\rangle=\sum_{i\in[d]}x_{i}y_{i}$ and write the resulting norm as $\|x\|=\sqrt{\langle x,x\rangle}$. Denote the set of unit vectors of $\mathbb{R}^{d}$ by $S^{d-1}$.
We endow the space of matrices $\mathbb{R}^{d\times d}$ with the standard operator norm. We write $M(d)=\mathbb{R}^{d\times d}$ and let $B_{M(d)}$ denote the unit ball in $M(d)$ with respect to the operator norm (i.e., the set of contractions).
Given a normed vector space $(V,\|\>\>\|)$ with $V\subseteq\mathbb{R}^{d}$, the dual norm of an element $v\in V$ is given by
$\|v\|_{*}=\sup\{|\langle v,w\rangle|\mid w\in V,\ \|w\|_{V}\leq 1\}.$
We write $\mathbf{i}$ for a $t$-tuple $\mathbf{i}=(i_{t},\dots,i_{t})\in[n]^{t}$ of indices. Given variables $x_{1},\ldots,x_{n}$ and a $t$-tuple $\mathbf{i}\in[n]^{t}$, we use $x(\mathbf{i})$ to denote the monomial $x_{i_{1}}x_{i_{2}}\cdots x_{i_{t}}$. Similarly, given a matrix-valued map $A\colon[n]\to\mathbb{R}^{d\times d}$, we write $A(\mathbf{i}):=A(i_{1})A(i_{2})\dots A(i_{t})$. We let $\mathcal{S}_{n}$ be the symmetric group on $n$ elements.
2.2. Polynomials
As usual we let $\mathbb{R}[x_{1},\dots,x_{n}]$ be the ring of $n$-variate polynomials with real coefficients, whose elements we write as
(4) $p(x)=\sum_{\alpha\in\mathbb{Z}_{\geq 0}^{n}}c_{\alpha}x^{\alpha},$
where $x^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$ and $c_{\alpha}\in\mathbb{R}$. We define the support of $p$ by
$\operatorname{supp}(p)=\{\alpha\in\mathbb{Z}_{\geq 0}^{n}\,\mid\,c_{\alpha}\neq 0\}.$
For $\alpha\in\mathbb{Z}_{\geq 0}^{n}$, write $|\alpha|=\alpha_{1}+\cdots+\alpha_{n}$, which is the degree of the monomial $x^{\alpha}$. A form of degree $d$ is a homogeneous polynomial of degree $d$, i.e., a polynomial whose support consists of $\alpha$ for which $|\alpha|=d$. Denote by $\mathbb{R}[x_{1},\dots,x_{n}]_{=d}$ the space of forms of degree $d$. For $p$ as in Eq. 4, define its homogeneous degree-$d$ part by
$p_{=d}(x)=\sum_{|\alpha|=d}c_{\alpha}x^{\alpha}.$
We endow $\mathbb{R}[x_{1},\dots,x_{n}]$ with the inner product given by
$\langle p,q\rangle=\sum_{\alpha\in\mathbb{Z}_{\geq 0}^{n}}c_{\alpha}c_{\alpha}^{\prime},$
where $c_{\alpha}$ and $c_{\alpha}^{\prime}$ are the coefficients of $p$ and $q$, respectively.
Multilinear polynomials (those that are affine in every variable) will play an important role in this work, as they can be identified with the real-valued functions defined on the Boolean hypercube.
2.3. Polynomials, tensors, norms and quantum query complexity
We recall the definition of $\|\cdot\|_{1}$ and $\|\cdot\|_{\infty}$, which are seminorms of polynomials in $\mathbb{R}[x_{1},\dots,x_{n}]$, and norms on the space of multilinear polynomials.
$\|p\|_{\infty}:=\sup_{x\in\{-1,1\}^{n}}|p(x)|,$
$\|p\|_{1}:=\mathbb{E}_{x\in\{-1,1\}^{n}}|p(x)|,$
where the expectation is taken with respect to the uniform probability measure.
Toward defining the completely bounded norm, we restrict our attention to forms only. For every $p\in\mathbb{R}[x_{1},\dots,x_{n}]_{=t}$ there are many $t$-tensors $T\in\mathbb{R}^{n\times\dots\times n}$ such that $T(x)=p(x)$ for every $x\in\mathbb{R}^{n}$, where
$T(x):=\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}x(\mathbf{i}).$
These tensors only have to satisfy
(5) $\sum_{\mathbf{i}\in\mathcal{I}_{\alpha}}T_{\mathbf{i}}=c_{\alpha}\ \ \ \forall\alpha\in\mathbb{Z}_{\geq 0}^{n},$
where $\mathcal{I}_{\alpha}$ is the set of elements $\mathbf{i}$ of $[n]^{t}$ such that the element $m\in[n]$ occurs $\alpha_{m}$ times in $\mathbf{i}$. Each of these tensors gives a way of evaluating $p$ in matrices, namely for every matrix-valued map $A:[n]\to M(d)$,
$T(A):=\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}A(\mathbf{i})=\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}A(i_{1})\dots A(i_{t}).$
We will use the following inner product for tensors:
$\langle T,R\rangle=\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}R_{\mathbf{i}}.$
It is also important to consider the unique symmetric $t$-tensor $T_{p}\in\mathbb{R}^{n\times\dots\times n}$ such that
$p(x)=T_{p}(x).$
For a tensor $T=(T_{\mathbf{i}})_{\mathbf{i}\in[n]^{t}}$ and permutation $\sigma\in\mathcal{S}_{t}$ we define $T\circ\sigma:=(T_{\sigma(\mathbf{i})})_{\mathbf{i}\in[n]^{t}}$, and we say that $T$ symmetric if $T=T\circ\sigma$ for all $\sigma\in\mathcal{S}_{t}$. The entries of this $T_{p}$ are given by
(6) $(T_{p})_{\mathbf{i}}=\frac{c_{e_{i_{1}}+\dots+e_{i_{t}}}}{\tau(i_{1},\dots,i_{t})}\ \mathrm{for}\ \mathbf{i}\in[n]^{t},$
where $\tau(i_{1},\dots,i_{t})$ is the number of distinct permutations of the sequence $(i_{1},\dots,i_{t})$. The completely bounded norm of a tensor $T$ is given by [^2] This is the completely bounded norm of $T$ when regarded as an element of $\ell^{1}_{n}\otimes_{h}\dots\otimes_{h}\ell^{1}_{n}$, where $h$ stands for the Haagerup tensor product, which determines a well-studied tensor norm. See for instance [Pau03, Chapter 17].
(7) $\|T\|_{{\mathrm{cb}}}=\sup$$\Big{\{}\Big{\|}\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}A_{1}(i_{1})\dots A_{t}(i_{t})\Big{\|}\,\mid\,A_{s}\colon[n]\to B_{M(d)},\ s\in[t],\ d\in\mathbb{N}\Big{\}}.$
It is not hard to show that for a $2$-tensor (matrix) $T$ this definition agrees with the one given in the introduction in Eq. 2.
The completely bounded norm of a form $p$ is given by
(8) $\|p\|_{{\mathrm{cb}}}=\inf\Big{\{}\sum_{\sigma\in\mathcal{S}_{t}}\|T^{\sigma}\|_{{\mathrm{cb}}}\,\mid\,\sum_{\sigma\in\mathcal{S}_{t}}T^{\sigma}\circ\sigma=T_{p}\Big{\}}.$
In Section 3.1 we provide easier expressions for both $\|T\|_{{\mathrm{cb}}}$ (showing that a single contraction-valued map $A\colon[n]\to B_{M(d)}$ suffices) and $\|p\|_{{\mathrm{cb}}}$. To prove our main technical result, Theorem 5.1, which expresses $\mathcal{E}(p,t)$ in the language of completely bounded polynomials, we combine these simplified expressions with the following characterization of quantum query algorithms, proved in [ABP19].
Theorem 2.1 (Quantum query algorithms are completely bounded forms).
Let $p:\{-1,1\}^{n}\to[-1,1]$ and let $t\in\mathbb{N}$. Then,
$\mathcal{E}(p,t)=\inf\quad$ $\|p-q\|_{\infty}$
$\mathrm{s.t.}\ \ \,$ $h\in\mathbb{R}[x_{1},\ldots,x_{n+1}]_{=2t}\text{ with }\|h\|_{{\mathrm{cb}}}\leq 1$
$q:\{-1,1\}^{n}\to\mathbb{R},\ \mathrm{with}\ q(x)=h(x,1)\ \forall\ x\in\{-1,1\}^{n}.$
2.4. Block-multilinear forms
Our main result Theorem 5.1 is stated for a special kind of polynomials, which are the block-multilinear forms.
Definition 2.2.
Let $\mathcal{P}=\{I_{1},\dots,I_{t}\}$ be a partition of $[n]$ into $t$ (pairwise disjoint) non-empty subsets. Define the set of block-multilinear polynomials with respect to $\mathcal{P}$ to be the linear subspace
$V_{\mathcal{P}}=\operatorname{Span}\big{\{}x_{i_{1}}\cdots x_{i_{t}}\,\mid\,i_{1}\in I_{1},\dots,i_{t}\in I_{t}\big{\}}.$
We also work with the larger space of polynomials spanned by monomials where in the above we replace linearity by odd degree.
Definition 2.3.
For a family $\mathcal{Q}\subseteq 2^{[m]}$ of pairwise disjoint subsets, let $W_{\mathcal{Q}}\subseteq\mathbb{R}[x_{1},\dots,x_{m}]$ be the subspace of polynomials spanned by monomials $x^{\alpha}$ with $\alpha\in\mathbb{Z}_{\geq 0}^{m}$ satisfying
(9) $\sum_{i\in I}\alpha_{i}\equiv 1\bmod 2\>\>\forall I\in\mathcal{Q}.$
We use $\Pi_{\mathcal{Q}}:\mathbb{R}[x_{1},\dots,x_{m}]\to W_{\mathcal{Q}}$ to refer to the projector onto $W_{\mathcal{Q}}.$
Remark 2.4.
Given a partition $\mathcal{P}$ of $[n]$, we have $V_{\mathcal{P}}\subset W_{\mathcal{P}}$. In particular, $V_{\mathcal{P}}$ consists of precisely the multilinear polynomials in $W_{\mathcal{P}}$.
Although the projector $\Pi_{\mathcal{Q}}$ onto $W_{\mathcal{Q}}$ is properly defined on the space of polynomials, we will slightly abuse notation and let it act on a $t$-tensor $T\in\mathbb{R}^{n\times\cdots\times n}$ as follows. Define $\mathcal{I}_{\mathcal{Q}}\subseteq[n]^{t}$ to be the set of $t$-tuples that contain an odd number of elements from each set $I\in\mathcal{Q}$. Then, we let $\Pi_{\mathcal{Q}}T$ be the tensor given by
(10) $(\Pi_{\mathcal{Q}}T)_{\mathbf{i}}:=\left\{\begin{array}[]{ll}T_{\mathbf{i}}&\mathrm{if}\ \mathbf{i}\in\mathcal{I}_{\mathcal{Q}},\\ 0&\mathrm{otherwise}.\end{array}\right.$
It is not hard to see that if $p$ is a polynomial satisfying $T(x)=p(x)$ for every $x\in\{-1,1\}^{n}$, then $\Pi_{\mathcal{Q}}T(x)=\Pi_{\mathcal{Q}}p(x)$ for every $x\in\{-1,1\}^{n}$.
We note that all the norms and seminorms we have mentioned are norms on the space $V_{\mathcal{P}}$ for any partition $\mathcal{P}$ of $[n]$. Hence, we can take the dual of these norms with respect to this subspace, so from now on $\|p\|_{\infty,*}$ and $\|p\|_{{\mathrm{cb}},*}$ will be the dual of $\|p\|_{\infty}$ and $\|p\|_{{\mathrm{cb}}}$ of $V_{\mathcal{P}}$, respectively. By contrast, when we write $\|R\|_{{\mathrm{cb}},*}$ for some $t$-tensor $\mathbb{R}^{n\times\dots\times n}$ we refer to the dual norm of the completely bounded norm of $R$ with respect to the whole space of $t$-tensors.
We stress that $*\|\cdot\|_{\infty,*}$ need not be equal to $\|\cdot\|_{1}$. This is because we are taking the dual norms with respect to $V_{\mathcal{P}}$ and not with respect to the space of all multilinear maps, in which case the dual norm would be $\|p\|_{1}$. The following example shows that $\|p\|_{\infty,*}\neq\|p\|_{1}$ in general.
Example 2.5.
Consider $n=3$, $t=1$ and $p=(x_{1}+x_{2}+x_{3})/3$. Then, $\|p\|_{1}>1/3$, but $\|p\|_{\infty,*}\leq 1/3$. Indeed, as $|p(x)|\geq 1/3$ for every $x\in\{-1,1\}^{3}$ and $|p(x)|>1/3$ for some $x\in\{-1,1\}^{3}$, we have that $\|p\|_{1}>1/3$. On the other hand, in this case $\mathcal{P}=\{[3]\}$ so $V_{\mathcal{P}}$ is the set of linear polynomials. Note that if $q$ is linear, then $\|\hat{q}\|_{1}=\|q\|_{\infty}$, where $\hat{q}$ is the Fourier transform of $q$. Hence
$\|p\|_{\infty,*}$ $=\sup_{q\in V_{\mathcal{P}},\|q\|_{\infty}\leq 1}\langle p,q\rangle=\sup_{q\in V_{\mathcal{P}},\|\hat{q}\|_{1}\leq 1}\langle\hat{p},\hat{q}\rangle\leq\sup_{\|\hat{q}\|_{1}\leq 1}\|\hat{p}\|_{\infty}\|\hat{q}\|_{1}=\frac{1}{3},$
where in second equality we used Parseval’s identity (see [O’D14, Chapter 1] for an introduction to Fourier analysis in the Boolean hypercube).
3. Theory of tensor and polynomial norms
3.1. Reformulation of the completely bounded norm of a form
In this section we prove a few general results regarding the completely bounded norms of tensors and polynomials. We provide simplifications of these norms that are not only helpful to develop our results, but might also facilitate the use of completely bounded polynomials to prove results in quantum query complexity. After this section, we will implicitly use the formulas of Propositions 3.1 and 3.2 as definitions for $\|p\|_{{\mathrm{cb}}}$ and $\|T\|_{{\mathrm{cb}}}$, respectively.
First, we provide a simpler expression (compared to Eq. 8) for $\|p\|_{{\mathrm{cb}}}$.
Proposition 3.1.
Let $p\in\mathbb{R}[x_{1},\dots,x_{n}]_{=t}$ be a form of degree $t$. Then,
$\|p\|_{{\mathrm{cb}}}=\inf\Big{\{}\|T\|_{{\mathrm{cb}}}\,\mid\,p(x)=T(x)\ \forall x\in\mathbb{R}^{n}\Big{\}}.$
  • Proof:
    We begin by noting that for any $t$-tensor $T$ and permutation $\sigma\in\mathcal{S}_{t}$, we have
    (11) $(T\circ\sigma)(x)$ $=\sum_{\mathbf{i}\in[n]^{t}}T_{\sigma(\mathbf{i})}x(\mathbf{i})=\sum_{\mathbf{i}\in[n]^{t}}T_{\sigma(\mathbf{i})}x(\sigma(\mathbf{i}))=T(x).$
    We first show that $\|p\|_{{\mathrm{cb}}}\leq\inf\{\|T\|_{{\mathrm{cb}}}\,\mid\,p(x)=T(x)\}$. By restricting in (8) to decompositions where $T^{\sigma}=T/t!$ for all $\sigma$, we have
    $\|p\|_{{\mathrm{cb}}}$ $=\inf\Big{\{}\sum_{\sigma\in\mathcal{S}_{t}}\|T^{\sigma}\|_{{\mathrm{cb}}}\,\mid\,T_{p}=\sum_{\sigma}T^{\sigma}\circ\sigma\Big{\}}$
    $\leq\inf\Big{\{}\|T\|_{{\mathrm{cb}}}\,\mid\,T_{p}=\frac{1}{t!}\sum_{\sigma}T\circ\sigma\Big{\}}.$
    Hence, it suffices to show that $T_{p}=\frac{1}{t!}\sum_{\sigma}T\circ\sigma$ if and only if $p(x)=T(x)$ for all $x\in\mathbb{R}^{n}$. For the ‘only if’ implication, assume $T_{p}=\frac{1}{t!}\sum_{\sigma}T\circ\sigma$, then for any $x\in\mathbb{R}^{n}$ we have that
    $p(x)=T_{p}(x)=\frac{1}{t!}\sum_{\sigma}T\circ\sigma(x)=\frac{1}{t!}\sum_{\sigma}T(x)=T(x),$
    where in the second equality we have used (11). For the reverse implication, assume that $p(x)=T(x)$ for all $x\in\mathbb{R}^{n}$. Then, using (11), we obtain
    $\frac{1}{t!}\sum_{\sigma}T\circ\sigma(x)=p(x)$
    for all $x\in\mathbb{R}^{n}$ and therefore $T_{p}=\frac{1}{t!}\sum_{\sigma}T\circ\sigma$ since $T_{p}$ is the unique symmetric $t$-tensor with $T_{p}(x)=p(x)$ for all $x\in\mathbb{R}^{n}$. This concludes the proof of the first inequality.
    Now for the other inequality, $\|p\|_{{\mathrm{cb}}}\geq\inf\{\|T\|_{{\mathrm{cb}}}\,\mid\,p(x)=T(x)\}$, let $T^{\sigma}$ ($\sigma\in\mathcal{S}_{t}$) be such that $T_{p}=\sum_{\sigma}T^{\sigma}\circ\sigma$ and define $T=\sum_{\sigma}T^{\sigma}$. Then we have
    $p(x)=T_{p}(x)=\sum_{\sigma}T^{\sigma}\circ\sigma(x)=\sum_{\sigma}T^{\sigma}(x)=T(x),$
    and $\|T\|_{{\mathrm{cb}}}\leq\sum_{\sigma}\|T^{\sigma}\|_{{\mathrm{cb}}}$ by the triangle inequality. $\Box$
Second, we show that the contraction-valued maps $A_{s}$ in the definition of $\|T\|_{{\mathrm{cb}}}$ (see Equation (7)) can be taken to be the same. This result can be understood as the fact that the polarization constant of completely bounded multilinear maps is $1$. [^3] In Banach space theory a $t$-linear map $T:X\times\dots\times X\to Y$ determines a homogeneous degree-$t$ polynomial $P:X\to Y:A\to T(A,\dots,A)$. The operator norms of $T$ and $P$ are equivalent if $T$ is symmetric: $\|T\|\leq\|P\|\leq K(t)\|T\|,$ where $K(t)$ is the polarization constant of of degree $t$. For a survey on the topic see [MMFPSS22, Section 5.1].
Proposition 3.2.
Let $T\in\mathbb{R}^{n\times\cdots\times n}$ be a $t$-tensor. Then,
$\|T\|_{{\mathrm{cb}}}=\sup\big{\{}\|T(A)\|\,\mid\,A\colon[n]\to B_{M(d)},\ d\in\mathbb{N}\big{\}}.$
  • Proof:
    Let ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|T\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ be the expression in the right-hand side of the statement. Note that it is the same as the expression of $\|T\|_{{\mathrm{cb}}}$, but now the contraction-valued maps $A_{1},\dots,A_{t}$ are all equal. This shows that ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|T\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\leq\|T\|_{{\mathrm{cb}}}$. To prove the other inequality, let $A_{1},\dots,A_{t}\colon[n]\to B_{M(d)}$ and $u,v\in S^{d-1}$. Now, define the contraction-valued map $A$ by $A(i):=\sum_{s\in[t]}e_{s}e_{s+1}^{\mathsf{T}}\otimes A_{s}(i)$ for $i\in[n]$, and define the unit vectors $u^{\prime}:=e_{1}\otimes u$ and $v^{\prime}:=e_{t+1}\otimes v.$ They satisfy
    $\langle u,A_{1}(i_{1})\dots A_{t}(i_{t})v\rangle=\langle u^{\prime},A(\mathbf{i})v^{\prime}\rangle\quad\text{for all }\mathbf{i}\in[n]^{t},$
    so in particular
    $\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}\langle u,A_{1}(i_{1})\dots A_{t}(i_{t})v\rangle=\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}\langle u^{\prime},A(\mathbf{i})v^{\prime}\rangle.$
    Taking the supremum over all maps $A_{s}$ and $u,v$ shows that $\|T\|_{{\mathrm{cb}}}\leq{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|T\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$, which concludes the proof. $\Box$
Third, we give a convenient formula for the dual of the completely bounded norm of a tensor. To state it in an compact way, we introduce the following subset of $t$-tensors in $\mathbb{R}^{n\times\cdots\times n}$:
(12) $\begin{split}K(n,t):=\{\langle u,A(\mathbf{i})v\rangle_{\mathbf{i}\in[n]^{t}}\,\mid\,d\in\mathbb{N},\ u,v\in S^{d-1},\ A\colon[n]\to B_{M(d)}\}.\end{split}$
Proposition 3.3.
Let $R\in\mathbb{R}^{n\times\dots\times n}$ be a $t$-tensor. Then,
(13) $\|R\|_{{\mathrm{cb}},*}=\inf\{w>0\,\mid\,R\in wK(n,t)\}.$
Remark 3.4.
Note that Proposition 3.3 states that $\|R\|_{{\mathrm{cb}},*}$ is the Minkowski norm defined by $K(n,t)$, so $K(n,t)$ is the unit ball of $(\mathbb{R}^{n\times\dots\times n},\|\cdot\|_{{\mathrm{cb}},*})$.
  • Proof:
[…]
    $\langle u,A(\mathbf{j})v\rangle=\begin{cases}1&\text{ if }\mathbf{j}=\mathbf{i},\\ 0&\text{ otherwise.}\end{cases}$
    To conclude that ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|R\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ is well-defined it then suffices to observe that the set of scalar multiples of elements in $K(n,t)$ is closed under addition. Indeed, if
    $R_{\mathbf{i}}=\langle u,A(\mathbf{i})v\rangle\ \mathrm{and}\ \tilde{R}_{\mathbf{i}}=\langle\tilde{u},\tilde{A}(\mathbf{i})\tilde{v}\rangle,$
    for some $u,v,\tilde{u},\tilde{v}\in\mathbb{R}^{d}$ with $\|u\|^{2}=\|v\|^{2}=w$, $\|\tilde{u}\|^{2}=\|\tilde{v}\|^{2}=\tilde{w}$ and maps $A,\tilde{A}\colon[n]\to B_{M(d)},$ then
    $R_{\mathbf{i}}+\tilde{R}_{\mathbf{i}}=\langle\hat{u},\hat{A}(\mathbf{i})\hat{v}\rangle,$
    where $\hat{u},\hat{v}\in\mathbb{R}^{2d}$ are the vectors with $\|\hat{u}\|^{2}=\|\hat{v}\|^{2}=w+\tilde{w}$ defined by $\hat{u}:=u\oplus\tilde{u}$, $v:=v\oplus\tilde{v}$ and the map $\hat{A}\colon[n]\to B_{M(2d)}$ is defined via
    $\hat{A}(i)=\begin{pmatrix}A(i)&0\\ 0&\tilde{A}(i)\end{pmatrix}.$
[…]
    Finally, note that the completely bounded norm of a $t$-tensor $R\in\mathbb{R}^{n\times\dots\times n}$ is given by
    $\|T\|_{{\mathrm{cb}}}=\sup\Big{\{}\big{|}\sum_{\mathbf{i}\in[n]^{t}}T_{\mathbf{i}}\langle u,A(\mathbf{i})v\rangle\big{|}\,\mid\,d\in\mathbb{N},\ u,v\in S^{d-1},\ A\colon[n]\to B_{M(d)}\Big{\}},$
    so $\|T\|_{{\mathrm{cb}}}={\left|\kern-1.07639pt\left|\kern-1.07639pt\left|T\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{*}$, and by the fact that the dual of the dual norm is the primal norm for finite-dimensional normed spaces, we conclude that
    ||||||=||||||**$*.{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}={\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{**}=\|\cdot\|_{{\mathrm{cb}},*}.$
    $\Box$
3.2. Completely bounded norms of bilinear forms
For a $t$-tensor $T$ and permutation $\sigma\in\mathcal{S}_{t}$, it is in general not true that $\|T\|_{{\mathrm{cb}}}=\|T\circ\sigma\|_{{\mathrm{cb}}}$ (see for instance [ABP19]). This is in part the reason for the complicated expression for the completely bounded norm of a form (8). It is not hard to show however, that when $T$ is a matrix we have $\|T^{\mathsf{T}}\|_{{\mathrm{cb}}}=\|T\|_{{\mathrm{cb}}}$. This gives the following reformulation of the completely bounded norm of forms of degree $2$.
Proposition 3.5.
Let $p\in\mathbb{R}[x_{1},\ldots,x_{n}]_{=2}$ be a quadratic form and let $T_{p}$ be the unique symmetric matrix associated to $p$ via (6). Then $\|p\|_{{\mathrm{cb}}}=\|T_{p}\|_{{\mathrm{cb}}}$.
  • Proof:
    Let $T\in\mathbb{R}^{n\times n}$ be a matrix. First, we have
    (14) $\|T^{\mathsf{T}}\|_{{\mathrm{cb}}}$ $=\sup\Big{\{}\|\sum_{i,j}T_{j,i}A(i)B(j)\|\,\mid\,A,B\colon[n]\to B_{M(d)}\Big{\}}$
    $=\sup\Big{\{}\|\sum_{i,j}T_{j,i}B(j)^{\mathsf{T}}A(i)^{\mathsf{T}}\|\,\mid\,A,B\colon[n]\to B_{M(d)}\Big{\}}$
    $=\|T\|_{{\mathrm{cb}}},$
    where we use (twice) that for any matrix $M$ we have $\|M\|=\|M^{\mathsf{T}}\|$.
    Let $T\in\mathbb{R}^{n\times n}$ be a matrix with $p(x)=T(x)$. Then, $T_{p}=(T+T^{\mathsf{T}})/2$, so from the above and the triangle inequality it follows that
    $\|T_{p}\|_{{\mathrm{cb}}}=\frac{1}{2}\|T+T^{\mathsf{T}}\|_{{\mathrm{cb}}}\leq\frac{1}{2}(\|T\|_{{\mathrm{cb}}}+\|T^{\mathsf{T}}\|_{{\mathrm{cb}}})=\|T\|_{{\mathrm{cb}}}.$
    Using Proposition 3.1 we conclude that $\|p\|_{{\mathrm{cb}}}=\|T_{p}\|_{{\mathrm{cb}}}$. $\Box$
Considering bilinear forms gives the following corollary.
Corollary 3.6.
Let $p:\{-1,1\}^{n}\times\{-1,1\}^{n}\to\mathbb{R}$ be a bilinear form and let $A\in\mathbb{R}^{n\times n}$ be such that $p(x,y)=x^{\mathsf{T}}Ay$ for all $x,y\in\mathbb{R}^{n}$. Then, $\|A\|_{{\mathrm{cb}}}=\|p\|_{{\mathrm{cb}}}$.
  • Proof:
    By Proposition 3.5 we have $\|p\|_{{\mathrm{cb}}}=\|T_{p}\|_{{\mathrm{cb}}}$. Now observe that $T_{p}=\frac{1}{2}\begin{pmatrix}0&A\\ A^{\mathsf{T}}&0\end{pmatrix}$ and hence
    $\|T_{p}\|_{{\mathrm{cb}}}$ $\leq\frac{1}{2}\left(\Big{\|}\begin{pmatrix}0&A\\ 0&0\end{pmatrix}\Big{\|}_{{\mathrm{cb}}}+\Big{\|}\begin{pmatrix}0&0\\ A^{\mathsf{T}}&0\end{pmatrix}\Big{\|}_{{\mathrm{cb}}}\right)$
    $\leq\frac{1}{2}\left(\|A\|_{{\mathrm{cb}}}+\|A^{\mathsf{T}}\|_{{\mathrm{cb}}}\right)$
    $=\|A\|_{{\mathrm{cb}}},$
    where the last equality uses (14).
    Conversely, let $B,C\colon[n]\to B_{M(d)}$. Note that for any unitary matrix $U$ we have $\|\sum_{i,j\in[n]}A_{i,j}B(i)C(j)\|=\|\sum_{i,j\in[n]}A_{i,j}B(i)C(j)U\|$. We may therefore assume without loss of generality that
    $\|\sum_{i,j\in[n]}A_{i,j}B(i)C(j)\|=\Big{|}\sum_{i,j\in[n]}A_{i,j}\langle u,B(i)C(j)u\rangle\Big{|},$
    for a single unit vector $u$. For every $i\in[n]$, define the following contractions:
    $Q_{1}(i)=B(i)=Q_{2}^{\mathsf{T}}(i),\qquad Q_{2}(n+i)=C(i)=Q_{1}^{\mathsf{T}}(n+i).$
    Then,
    $\|T_{p}\|_{{\mathrm{cb}}}$ $\geq\Big{\|}\sum_{i,j\in[n]}\frac{1}{2}\big{(}A_{i,j}Q_{1}(i)Q_{2}(n+j)+A_{j,i}Q_{1}(n+i)Q_{2}(j)\big{)}\Big{\|}$
    $=\Big{|}\sum_{i,j\in[n]}A_{i,j}\langle u,\frac{B(i)C(j)+C^{\mathsf{T}}(j)B^{\mathsf{T}}(i)}{2}u\rangle\Big{|}$
    $=\Big{|}\sum_{i,j\in[n]}A_{i,j}\langle u,B(i)C(j)u\rangle\Big{|}$
    Taking the supremum over $B,C\colon[n]\to B_{M(d)}$ shows that $\|T_{p}\|_{{\mathrm{cb}}}\geq\|A\|_{{\mathrm{cb}}}$. $\Box$
3.3. Contractivity of the projector $\Pi_{Q}$.
The map $\Pi_{\mathcal{Q}}$ introduced in Definition 2.3 turns out to have a particularly nice structure in the form of an averaging operator, which we will use to show that it is contractive in several norms that are relevant for us. Throughout this section, let $\mathcal{Q}$ be a family of disjoint subsets of $[n]$. For each $I\in\mathcal{Q}$ let $z_{I}$ be a random variable that takes the values $-1$ and $1$ with probability $1/2$ and let $z=(z_{I})_{I\in\mathcal{Q}}$. For a bit string $x\in\{-1,1\}^{n}$, we define the random variable $x\cdot z\in\{-1,1\}^{n}$ as
$(x\cdot z)(i):=\left\{\begin{array}[]{ll}x_{i}z_{I}&\mathrm{if\ }i\in I\ \mathrm{for\ some\ }I\in\mathcal{Q},\\ x_{i}&\mathrm{otherwise}.\end{array}\right.$
For a matrix-valued map $A:[n]\to M(d)$ we define the random variable $A\cdot z$ in an analogous way.
Proposition 3.7.
For any $p\in\mathbb{R}[x_{1},\dots,x_{n}]$ and $x\in\mathbb{R}^{n}$, we have that
$\Pi_{\mathcal{Q}}p(x)$ $=\mathbb{E}_{z}\Big{[}p(x\cdot z)\prod_{I\in\mathcal{Q}}z_{I}\Big{]}.$
Similarly, for any $t$-tensor $T\in\mathbb{R}^{n\times\dots\times n}$, positive integer $d$ and matrix-valued map $A\colon[n]\to M(d)$, we have that
$\Pi_{\mathcal{Q}}T(A)$ $=\mathbb{E}_{z}\left[T(A\cdot z)\prod_{I\in\mathcal{Q}}z_{I}\right].$
  • Proof:
    By linearity, it suffices to prove the equality for monomials. Let $\alpha\in\mathbb{Z}_{\geq 0}^{n}$. Then we have
    $(x\cdot z)^{\alpha}\prod_{I\in\mathcal{Q}}z_{I}=x^{\alpha}\prod_{I\in\mathcal{Q}}z_{I}^{1+\sum_{i\in I}\alpha_{i}}.$
    It follows that
    $\mathbb{E}_{z}\Big{[}(x\cdot z)^{\alpha}\prod_{I\in\mathcal{Q}}z_{I}\Big{]}=\begin{cases}x^{\alpha}&\text{if }1+\sum_{i\in I}\alpha_{i}=0\bmod 2\ \forall I\in\mathcal{Q},\\ 0&\text{otherwise}.\end{cases}$
    It remains to observe that this is precisely the projection of $x^{\alpha}$ on $W_{\mathcal{Q}}$. The statement for tensors follows analogously. $\Box$
We finally prove that $\Pi_{\mathcal{Q}}$ is contractive with respect to the norms that we care about.
Lemma 3.8.
Let $\mathcal{Q}$ be a family of disjoint subsets of $[n]$ and $p\in\mathbb{R}[x_{1},\dots,x_{n}]$ and let $\mathrm{norm}\in\{{\mathrm{cb}},\infty,1\}$ where for the ${\mathrm{cb}}$-norm we moreover require $p$ to be homogeneous. Then
$\|\Pi_{\mathcal{Q}}p\|_{\mathrm{norm}}\leq\|p\|_{\mathrm{norm}}.$
  • Proof:
    First, we consider the $\|\cdot\|_{\infty}$ norm. For every $x\in\{-1,1\}^{n}$, we have that $x\cdot z\in\{-1,1\}^{n}$, so
    $|\Pi_{\mathcal{Q}}p(x)|\leq\mathbb{E}_{z}|p(x\cdot z)\prod_{I\in\mathcal{Q}}z_{I}|=\mathbb{E}_{z}|p(x\cdot z)|\leq\mathbb{E}_{z}\|p\|_{\infty}=\|p\|_{\infty},$
    where in the first inequality we used Proposition 3.7 and the triangle inequality.
    Second, we consider $\|\cdot\|_{\mathrm{cb}}$. Arguing as in the $\|\cdot\|_{\infty}$ case and using Proposition 3.2, it follows that for any $t$-tensor $T\in\mathbb{R}^{n\times\dots\times n}$ we have that $\|\Pi_{\mathcal{Q}}T\|_{{\mathrm{cb}}}\leq\|T\|_{{\mathrm{cb}}}$. Given that $\Pi_{\mathcal{Q}}p(x)=\Pi_{\mathcal{Q}}T(x)$ if $p(x)=T(x)$, it follows that
    $\|\Pi_{\mathcal{Q}}p\|_{{\mathrm{cb}}}\leq\|\Pi_{\mathcal{Q}}T\|_{{\mathrm{cb}}}\leq\|T\|_{{\mathrm{cb}}}$
    for every $t$-tensor $T\in\mathbb{R}^{n\times\dots\times n}$ such that $T(x)=p(x)$. Taking the infimum over all those $T$ we arrive at $\|\Pi_{\mathcal{Q}}p\|_{{\mathrm{cb}}}\leq\|p\|_{{\mathrm{cb}}}$.
    Finally, for $\|\cdot\|_{1}$ we have
    $\|\Pi_{\mathcal{Q}}p\|_{1}=\mathbb{E}_{x}|\mathbb{E}_{z}p(x\cdot z)\prod_{I\in\mathcal{Q}}z_{I}|\leq\mathbb{E}_{x}\mathbb{E}_{z}|p(x\cdot z)|=\mathbb{E}_{z}\mathbb{E}_{x}|p(x)|=\|p\|_{1},$
    where in the first equality we have used Proposition 3.7 and in the third we have used the fact that the uniform measure is invariant under products of $z\in\{-1,1\}^{n}$. $\Box$
4. No multiplicative converse for the polynomial method
In this section we prove Theorem 1.3, which rules out a multiplicative converse of the polynomial method. For this, we use the following straightforward corollary of Theorem 2.1.
Corollary 4.1.
Let $p:\{-1,1\}^{n}\to[-1,1]$ and let $t\in\mathbb{N}$. Suppose $\mathcal{A}$ is a $t$-query quantum algorithm such that $\mathbb{E}[\mathcal{A}(x)]=p(x)$ for every $x\in\{-1,1\}^{n}.$ Then, there exists a polynomial $h\in\mathbb{R}[x_{1},\dots,x_{n+1}]$ such that $h(x,1)=p(x)$ for every $x\in\{-1,1\}^{n}$ and $\|h\|_{\mathrm{cb}}\leq 1$.
This corollary shows that it suffices to exhibit a bounded quartic polynomial $p$ (or rather, an infinite sequence of such polynomials) such that any extension $h$ by one extra variable has unbounded completely bounded norm.
First, we will give a general way of lower bounding the completely bounded norm of such extensions. Then, we will explain what was the error in the original proof of [ABP19]. Finally, we will give two sequences of quartic polynomials that meet the requirements.
4.1. Lower bounding the completely bounded norm
In this subsection we will give a technique to lower bound the completely bounded norm of extensions of multilinear quartic forms (Corollary 4.3). The main technical tool we use is Lemma 4.2, a slight variant of a decomposition due to Varopoulos [Var74].
We will first talk about general cubic forms, that is polynomials given by:
(15) $p(x)=\sum_{S\in{[n]\choose 3}}c_{S}\prod_{i\in S}x_{i},$
where the $c_{S}$ are some real coefficients. We will lower bound its completely bounded norm. Then, we will extent this lower bound to an associated quartic form, given by $x_{0}p(x)$. For $i\in[n]$, define the $i$th slice of $p$ to be the symmetric matrix $M_{i}\in\mathbb{R}^{n\times n}$ with $(j,k)$-coefficient equal to $c_{\{i,j,k\}}$ if $i,j,k$ are pairwise distinct and 00 otherwise. Then, define
$\Delta(p)=\max_{i\in[n]}\|M_{i}\|.$
Lemma 4.2 (tri-linear Varopoulos decomposition).
Let $p$ be an $n$-variate multilinear cubic form as in (15). Then, for some $d\in\mathbb{N}$, there exists  $A:[n]\to B_{M(d)}$ and orthogonal vectors $u,v\in S^{d-1}$ such that $[A(j),A(i)]=0$, and
(16) $A(i)^{2}$ $=0$
(17) $\langle u,A(i)v\rangle$ $=0$
(18) $\langle u,A(i)A(j)v\rangle$ $=0$
(19) $\langle u,A(i)A(j)A(k)v\rangle$ $=\frac{c_{\{i,j,k\}}}{\Delta(p)}$
for all pairwise distinct $i,j,k\in[n]$.
  • Proof:
    For each $i\in[n]$, define $M_{i}$ as above. Define $W_{i}=\Delta(p)^{-1}M_{i}$ and note that $W_{i}$ has operator norm at most 1. For each $i\in[n]$, define the $(2n+2)\times(2n+2)$ block matrix
    $A(i)={\footnotesize\left[\begin{array}[]{c|c|c|c}&&&\\ \hline\cr e_{i}&&&\\ \hline\cr&W_{i}^{\mathsf{T}}&&\\ \hline\cr&&e_{i}^{\mathsf{T}}&\end{array}\right]},$
    where the first and last rows and columns have size 1, the second and third have size $n$ and where the empty blocks are filled with zeros. Define $u=e_{2n+1}$ and $v=e_{1}$. The rest of the proof is identical to the proof of [BP19, Lemma 2.11], except for the property that $A(i)^{2}=0$. This follows from the fact that
    $A(i)^{2}={\footnotesize\left[\begin{array}[]{c|c|c|c}&&&\\ \hline\cr&&&\\ \hline\cr W_{i}^{\mathsf{T}}e_{i}&&&\\ \hline\cr&e_{i}^{\mathsf{T}}W_{i}^{\mathsf{T}}&&\end{array}\right]}$
    and that the $i$th row and $i$th column of $M_{i}$ (and hence $W_{i}$) are zero. $\Box$
Corollary 4.3.
Let $p$ be an $n$-variate multilinear cubic form as in (15). Suppose that an $(n+2)$-variate quartic form $h\in\mathbb{R}[x_{0},x_{1},\dots,x_{n},z]$ satisfies $h(x,1)=x_{0}p(x_{1},\dots,x_{n})$ for every $x\in\{-1,1\}^{n+1}$. Then,
$\|h\|_{{\mathrm{cb}}}\geq\frac{\|p\|_{2}^{2}}{\Delta(p)}.$
  • Proof:
    From the orthonormality of the characters, it follows that $h$ and $x_{0}p$ have equal coefficients for each quartic multilinear monomial in the variables $x_{0},\dots,x_{n}$, which are $c_{S}$ for $x_{0}\chi_{S}$ with $S\in{[n]\choose 3}$ and 0 otherwise. Let $A:[n]\to B_{M(d)}$ and $u,v\in S^{d}$ be as in Lemma 4.2, and extend $A$ by $A(0)=I,A(n+1)=0$. Commutativity and properties (16)–(18) imply that if a quartic monomial expression $A((i,j,k,l))$ with $i,j,k,l\in\{0,\dots,n+1\}$ has repeated indices or an index equal to $n+1$, then $\langle u,A((i,j,k,l))v\rangle=0$. With this, it follows that, for every $T_{h}$ such that $T_{h}(x)=h(x)$, we have
    (20) $\|T_{h}\|_{{\mathrm{cb}}}$ $\geq\sum_{\mathbf{i}\in(\{0\}\cup[n+1])^{4}}T_{\mathbf{i}}\Big{\langle}u,A(\mathbf{i})v\Big{\rangle}=\sum_{S\in{[n]\choose 3}}c_{S}\Big{\langle}u,\,A(0)\prod_{i\in S}A(i)v\Big{\rangle}.$
    Finally, if we use that $A(0)=I$, property (19) and Parseval’s identity, we obtain the desired result:
    $\|h\|_{{\mathrm{cb}}}=\inf\|T_{h}\|_{{\mathrm{cb}}}$ $\geq\sum_{S\in{[n]\choose 3}}c_{S}\langle u,\prod_{i\in S}A(i)v\rangle=\Delta(p)^{-1}\sum_{S\in{[n]\choose 3}}c_{S}^{2}=\frac{\|p\|_{2}^{2}}{\Delta(p)}.$
    $\Box$
4.2. Error in the original proof of Theorem 1.3.
Here, we discuss the error in [ABP19, pp. 920]. Their proof arrives at the equation
(21) $\sum_{\alpha,\beta\in\{0,1,2,3,4\}^{n}:|\alpha|+|\beta|=4}d^{\prime}_{\alpha,\beta}x^{\alpha}=C\sum_{\alpha\in\{0,1\}^{n}:|\alpha|=4}d_{\alpha}x^{\alpha}\ \ \forall\ x\in\{-1,1\}^{n},$
where $d^{\prime}_{\alpha,\beta}$, $d_{\alpha}$ and $C$ are some real numbers, $x^{\alpha}$ stands for $\prod_{i=1}^{n}x_{i}^{\alpha_{i}}$ and $|\alpha|$ for $\sum_{i=1}^{n}\alpha_{i}$. It follows from the orthogonality of the characters that $d^{\prime}_{\alpha,0}=Cd_{\alpha}$ for all $\alpha\in\{0,1\}^{n}$ such that $|\alpha|=4$. What is used, however, is that $d^{\prime}_{\alpha,0}=Cd_{\alpha}$ for all $\alpha\in\{0,1,2,3,4\}^{n}$ such that $|\alpha|=4$, which is not true in general. For instance if $n=1$, $C=1$ and $d^{\prime}_{(2,0),(0,2)}=-d^{\prime}_{(0,0),(4,0)}=1$ and the rest of the coefficients set to 00, then (21) becomes $x^{2}-1=0,\ \forall\ x\in\{-1,1\}$.
Corollary 4.3 gets around this issue by using a multilinear cubic form instead of just a cubic form. This results in matrices $A(i)$ in Lemma 4.2 that square to zero and has the effect that terms other than quartic multilinear monomials vanish in the left-hand side of (20).
4.3. A proof based on a random example
The probabilistic proof of Theorem 1.3 uses a random cubic form as in (15) where the coefficients $c_{S}$ are chosen to be independent uniformly distributed random signs. Parseval’s identity then gives $\|p\|_{2}^{2}={n\choose 3}$. Each of the slices $M_{i}$ of $p$ is a random symmetric matrix with independent mean-zero entries of absolute value at most 1. A standard random-matrix inequality and the union bound then imply that $\Delta(p)\leq C\sqrt{n}$ with probability $1-\exp(-Cn)$ [Tao12, Corollary 2.3.6]. By Hoeffding’s inequality [BLM13, Theorem 2.8] and the union bound, we have that $\|p\|_{\infty}\leq Cn^{2}$ with probability $1-\exp(-Cn)$. Rescaling $p$ then gives that there exists a bounded multilinear cubic form such that $\|p\|_{2}^{2}/\Delta(p)\geq C\sqrt{n}$.
Now, if Theorem 1.3 was false, by Corollary 4.1, there would be a constant $C^{\prime}\in(0,1]$ and $h:\{-1,1\}^{n+2}\to\mathbb{R}$ such that $h(x,1)=C^{\prime}x_{0}p(x_{1},\dots,x_{n})$ and $1\geq\|h\|_{\mathrm{cb}}$. On the other hand, Corollary 4.3 implies that $\|h\|_{\mathrm{cb}}\geq CC^{\prime}\sqrt{n}$. Putting both things together, we reach a contradiction.
4.4. A proof based on an explicit example
We also give a constructive proof of Theorem 1.3 using techniques from [BP19], which were used there to disprove a conjecture on a tri-linear version of Grothendieck’s theorem. We do not exactly use the construction from that paper because it involves complex functions. Instead, we will use the Möbius function (defined below), which is real valued and has the desired properties.
The construction uses some definitions from additive combinatorics. For functions $f_{1},f_{2},f_{3},f_{4}:\mathbb{Z}_{n}\to[-1,1]$ (on the cyclic group of order $n$), define the 4-linear form
$\Lambda_{4}(f_{1},f_{2},f_{3},f_{4})={\mathbb{E}}_{a,b\in\mathbb{Z}_{n}}f_{1}(a)f_{2}(a+b)f_{3}(a+2b)f_{4}(a+3b),$
where the expectation is taken with respect to independent and uniformly distributed $a,b\in\mathbb{Z}_{n}$. For a function $f_{4}:\mathbb{Z}_{n}\to[-1,1]$ to be set later, we will use this to define a tri-linear form on $3n$ variables as
$p(x)=\sum_{a,b\in\mathbb{Z}_{n}}x_{1,a}x_{2,a+b}x_{3,a+2b}f_{4}(a+3b),$
where the variables are indexed by $[3]\times\mathbb{Z}_{n}$. Using the identification $f_{i}(a)=x_{i,a}$ for $i\in[3]$, we find the relation $p=n^{2}\Lambda_{4}$. In particular,
(22) $\|p\|_{\infty}=n^{2}\max_{f_{1},f_{2},f_{3}:\mathbb{Z}_{n}\to[-1,1]}|\Lambda_{4}(f_{1},f_{2},f_{3},f_{4})|.$
This norm can be bounded from above in terms of the Gowers 3-uniformity norm, which for a function $f:\mathbb{Z}_{n}\to[-1,1]$ is defined by
$\|f\|_{U^{3}}=\Big{(}{\mathbb{E}}_{a,b_{1},b_{2},b_{3}\in\mathbb{Z}_{n}}\prod_{c\in\{0,1\}^{3}}f(a+c_{1}b_{1}+c_{2}b_{2}+c_{3}b_{3})\Big{)}^{\frac{1}{8}}.$
We then have the following basic bound [Gre07, Proposition 1.11].
Lemma 4.4 (generalized von Neumann inequality).
Suppose that $n$ is coprime to 6. Then, for any $f_{1},f_{2},f_{3},f_{4}:\mathbb{Z}_{n}\to[-1,1]$, we have that
$|\Lambda_{4}(f_{1},f_{2},f_{3},f_{4})|\leq\min_{i\in[4]}\|f_{i}\|_{U^{3}}.$
Combining this lemma with (22), it follows that
$\|p\|_{\infty}\leq n^{2}\|f_{4}\|_{U^{3}}.$
The polynomial $p$ has $3n$ slices, $M_{i,a}\in\mathbb{R}^{[3]\times\mathbb{Z}_{n}}$ for each $i\in[3]$ and $a\in\mathbb{Z}_{n}$, which we view as $3\times 3$ block-matrices with blocks indexed by $\mathbb{Z}_{n}$. The slice $M_{1,a}$ is supported only on the $(2,3)$ and $(3,2)$ blocks, which are each others’ transposes. On its $(2,3)$ block it has value $f_{4}(a+3b)$ on coordinate $(a+b,a+2b)$ for each $b$. In particular, this matrix has at most one nonzero entry in each row and column. It follows that a relabeling of the rows turns $M_{1,a}$ into a diagonal matrix with diagonal entries in $[-1,1]$, and therefore $\|M_{1,a}\|\leq 1$. Similarly, we get that $\|M_{i,a}\|\leq 1$ for $i=2,3$. Hence, $\Delta(p)\leq 1$. Parseval’s identity implies that
$\|p\|_{2}^{2}=n\sum_{a\in\mathbb{Z}_{n}}f_{4}(a)^{2}.$
Identify $\mathbb{Z}_{n}$ with $\{0,1\dots,n-1\}$ in the standard way. We choose $f_{4}$ to be the Möbius function restricted to this interval. That is, set $f_{4}(0)=0$ and for $a>0$, set
$f_{4}(a)=\begin{cases}1&\text{if~{}$a$ is square-free with an even number of prime factors}\\ -1&\text{if~{}$a$ is square-free with an odd number of prime factors}\\ 0&\text{otherwise.}\end{cases}$
Tao and Teräväinen [TT21] recently proved that
$\|f_{4}\|_{U^{3}}\leq\frac{1}{(\log\log n)^{C}}$
for some absolute constant $C>0$. Since there are $\frac{6}{\pi^{2}}n-O(\sqrt{n})$ integers in $[n]$ that are square-free [${}^{+}$79, page 269], normalizing $p$ by $(\log\log n)^{C}/n^{2}$ and taking $n$ coprime to 6 gives a bounded multilinear cubic polynomial satisfying
$\frac{\|p\|_{2}^{2}}{\Delta(p)}\geq\frac{6}{\pi^{2}}(\log\log n)^{C}-o(1).$
This proves Theorem 1.3 as before.
Remark 4.5.
The jointly completely bounded norm of $p$ is given by
$\|p\|_{\operatorname{jcb}}=\sup_{d\in\mathbb{N}}\|\sum_{a,b\in\mathbb{Z}_{n}}A(1,a)A(2,a+b)A(3,a+2b)f_{4}(a+3b)\|,$
where the supremum is taken over maps $A:[3]\times[n]\to\mathbb{C}^{d\times d}$ such that $\|A(i,a)\|\leq 1$ and $[A(i,a),A(j,b)]=[A(i,a),A(j,b)^{*}]=0$ for all ${i\neq j}$ and $a,b\in\mathbb{Z}_{n}$. This norm can also be stated in terms of tensor products and the supremum is attained by observable-valued maps. As such, this norm appears naturally in the context of non-local games. It was shown in [${}^{+}$19] that Proposition 4.4 also holds for the jointly completely bounded norm, that is $\|p\|_{\operatorname{jcb}}\leq n^{2}\|f_{4}\|_{U^{3}}$. The proof of Corollary 4.3 easily implies that $\|p\|_{{\mathrm{cb}}}\geq\|p\|_{2}^{2}/\Delta(p)$. This was used in [BP19] to prove that the $\operatorname{jcb}$ and ${\mathrm{cb}}$ norms are inequivalent.
5. $\mathcal{E}(p,t)$ for block-multilinear forms
In this section we will give two proofs of Theorem 5.1 and also provide formulas to compute the dual norms appearing in our results.
5.1. First proof of main result
First of all, we formally state our main result Theorem 5.1.
Theorem 5.1.
Let $\mathcal{P}$ be a partition of $[n]$ in $2t$ subsets and $p\in V_{\mathcal{P}}$. Then,
$\mathcal{E}(p,t)=\sup\left\{\langle p,r\rangle-\|r\|_{{\mathrm{cb}},*}\,\mid\,r\in V_{\mathcal{P}},\ \|r\|_{\infty,*}\leq 1\right\}.$
  • Proof:
    We start from the expression given in Theorem 2.1 for $\mathcal{E}(p,t)$ and let $h\in\mathbb{R}[x_{1},\ldots,x_{n+1}]_{=2t}\text{ with }\|h\|_{{\mathrm{cb}}}\leq 1$ and let $q:\{-1,1\}^{n}\to\mathbb{R}$ be defined by $q(x)=h(x,1)$ for every $x\in\{-1,1\}^{n}$. We first show that we can project $q$ (and $h$) onto $W_{\mathcal{P}}$ and obtain a feasible solution whose objective value is at least as good as $q$. Since $\mathcal{P}$ is a partition of $[n]$, it defines a family of disjoint subsets of $[n+1]$, so by Lemma 3.8, we have $\|\Pi_{\mathcal{P}}h\|_{{\mathrm{cb}}}\leq\|h\|_{{\mathrm{cb}}}\leq 1$. Since the degree of $h$ is at most $2t$, the polynomial $\Pi_{\mathcal{P}}h$ has degree at most $2t$. This shows that each monomial in its support contains exactly one variable from each of the $2t$ sets in $\mathcal{P}$. We can therefore observe that $\Pi_{\mathcal{P}}h$ does not depend on $x_{n+1}$. Since $h(x,1)=q(x)$, we have $\Pi_{\mathcal{P}}h(x,1)=\Pi_{\mathcal{P}}q(x)$ and therefore $\Pi_{\mathcal{P}}q\in V_{\mathcal{P}}$. We then use Proposition 3.1 to show that $\|\Pi_{\mathcal{P}}q\|_{{\mathrm{cb}}}\leq 1$. Indeed, applying $\Pi_{\mathcal{P}}$ to a $2t$-tensor $T\in\mathbb{R}^{(n+1)\times\cdots(n+1)}$ that certifies $\|h\|_{{\mathrm{cb}}}\leq 1$ results in a tensor $\Pi_{\mathcal{P}}T$ that satisfies $\Pi_{\mathcal{P}}T(\mathbf{i})=0$ whenever $\mathbf{i}$ contains an index equal to $n+1$. So, $\Pi_{\mathcal{P}}T(x,1)=\Pi_{\mathcal{P}}q(x)$ for every $x\in\{-1,1\}^{n}$ and thus $\Pi_{\mathcal{P}}T$, viewed as a $2t$-tensor in $\mathbb{R}^{n\times\cdots\times n}$, certifies $\|\Pi_{\mathcal{P}}q\|_{{\mathrm{cb}}}\leq 1$. For the objective value of $\Pi_{\mathcal{P}}q$ we finally observe that
    $\|p-\Pi_{\mathcal{P}}q\|_{\infty}=\|\Pi_{\mathcal{P}}(p-q)\|_{\infty}\leq\|p-q\|_{\infty},$
    where we used that $p\in V_{\mathcal{P}}$ in the equality and Lemma 3.8 in the inequality. This shows that
    $\mathcal{E}(p,t)\geq\inf\{\|p-q\|_{\infty}\,\mid\,q\in V_{\mathcal{P}}\text{ with }\|q\|_{{\mathrm{cb}}}\leq 1\}.$
    To show that the above inequality is in fact an equality it suffices to observe that given a polynomial $q\in V_{\mathcal{P}}$, we can define $h\in\mathbb{R}[x_{1},\ldots,x_{n+1}]$ as $h(x,x_{n+1})=q(x)$ and then we have $\|h\|_{{\mathrm{cb}}}\leq\|q\|_{{\mathrm{cb}}}$. Finally, in the above reformulation of $\mathcal{E}(p,t)$, we can express $\|p-q\|_{\infty}$ in terms of its dual norm and obtain
    $\mathcal{E}(p,t)=\inf_{q}\sup_{r}\quad$ $\langle p-q,r\rangle$
    $\mathrm{s.t.}\quad$ $q\in V_{\mathcal{P}}\ \mathrm{with\ }\|q\|_{{\mathrm{cb}}}\leq 1,$
    $r\in V_{\mathcal{P}}\ \mathrm{with\ }\|r\|_{\infty,*}\leq 1.$
    The desired result then follows by exchanging the infimum and supremum, which we are allowed to do by von Neumann’s minimax theorem (see, e.g., [Nik54] for a simple statement and proof) and the definition of $\|r\|_{{\mathrm{cb}},*}$. $\Box$
5.2. How to compute the dual norms
Theorem 5.1 provides a formula for $\mathcal{E}(p,t)$ when $p\in V_{\mathcal{P}}$ that is easier to work with than the one of Theorem 2.1, as we will see in the next sections. However, it is not clear yet how to compute the dual norms appearing in Theorem 5.1. Now, we will give expressions for both norms.
Proposition 5.2.
Let $\mathcal{P}$ be a partition of $[n]$ in $t$ subsets and $p\in V_{\mathcal{P}}$. Then,
(23) $\|p\|_{\infty,*}=\inf\{\|r\|_{1}\,\mid\,\,r:\{-1,1\}^{n}\to\mathbb{R},\,r\in W_{\mathcal{P}},\,r_{=t}=p\big{\}}.$
  • Proof:
    First, by Lagrange duality (cf. [BV04, Sec. 5.1.6]), we have
    (24) $\inf\big{\{}\|r\|_{1}\,\mid\,r:\{-1,1\}^{n}\to\mathbb{R},\,r_{=t}=p\big{\}}$
    $=\sup\big{\{}\langle p,q\rangle\,\mid\,q\in\mathbb{R}[x_{1},\dots,x_{n}]_{=t},\,\|q\|_{\infty}\leq 1\}$
    $=\sup\big{\{}\langle p,q\rangle\,\mid\,q\in W_{\mathcal{P}},\,\|q\|_{\infty}\leq 1\}=\|p\|_{\infty,*},$
    where for the second equality we use that $p\in V_{\mathcal{P}}$ and we replace $q$ by $\Pi_{\mathcal{P}}q$ (which does not increase its infinity-norm by Lemma 3.8). Second, we show that the right-hand side of Eq. 23 equals (24). The lower bound follows from inclusion of the feasible region, while the upper bound follows from Lemma 3.8 for $\|\cdot\|_{1}$. $\Box$
We have already seen in Example 2.5 that $\|p\|_{\infty,*}\neq\|p\|_{1}$ in general, because we are taking the dual norm with respect to $V_{\mathcal{P}}$. For completeness we give an alternative proof of the separation using Proposition 5.2.
Example 5.3.
The upper bound $\|p\|_{\infty,*}\leq 1/3$ of Example 2.5 follows from Proposition 5.2 by considering the multilinear map $r(x)=(x_{1}+x_{2}+x_{3}+x_{1}x_{2}x_{3})/3$ that belongs to $W_{\mathcal{P}}$, satisfies $r_{=1}(x)=p(x)=(x_{1}+x_{2}+x_{3})/3$ and $\|r\|_{1}=1/3$.
Proposition 5.4.
Let $\mathcal{P}$ be a partition of $[n]$ in $t$ subsets and let $p\in V_{\mathcal{P}}$. Then,
(25) $\|p\|_{{\mathrm{cb}},*}=t!\|T_{p}\|_{{\mathrm{cb}},*}.$
  • Proof:
    By duality and definition of $\|\cdot\|_{{\mathrm{cb}}}$, we have that
    $\|p\|_{{\mathrm{cb}},*}$ $=\sup\{\sum_{\alpha\in\mathbb{Z}_{\geq 0}^{n}}c_{\alpha}c_{\alpha}^{\prime}:q\in V_{\mathcal{P}},\,\|q\|_{{\mathrm{cb}}}\leq 1\}$
    $=\sup\{\sum_{\alpha\in\mathbb{Z}_{\geq 0}^{n}}c_{\alpha}\sum_{\mathbf{i}\in\mathcal{I}_{\alpha}}T_{\mathbf{i}}:\,q\in V_{\mathcal{P}},\,T(x)=q(x),\,\|T\|_{{\mathrm{cb}}}\leq 1\},$
    where $c_{\alpha}$ and $c_{\alpha}^{\prime}$ are the coefficients of $p$ and $q$, respectively. Now, let $R\in\mathbb{R}^{n\times\dots\times n}$ be a $t$-tensor such that $R(x)=p(x)$ for every $x\in\mathbb{R}^{n}$. Then we have, using Eq. 5, that
    $\|p\|_{{\mathrm{cb}},*}$ $=\sup\{\sum_{\alpha\in\mathbb{Z}_{\geq 0}^{n}}\sum_{\mathbf{j}\in\mathcal{I}_{\alpha}}R_{\mathbf{j}}\sum_{\mathbf{i}\in\mathcal{I}_{\alpha}}T_{\mathbf{i}}:q\in V_{\mathcal{P}},\,T(x)=q(x),\,\|T\|_{{\mathrm{cb}}}\leq 1\}.$
    In particular, if we choose $R$ to be $T_{p}$, then we have
    $\|p\|_{{\mathrm{cb}},*}$ $=t!\sup\{\sum_{\alpha\in\mathbb{Z}_{\geq 0}^{n}}\sum_{\mathbf{i}\in\mathcal{I}_{\alpha}}(T_{p})_{\mathbf{i}}T_{\mathbf{i}}:q\in V_{\mathcal{P}},\,T(x)=q(x),\,\|T\|_{{\mathrm{cb}}}\leq 1\}$
    $=t!\sup\{\langle T_{p},T\rangle:q\in V_{\mathcal{P}},\,T(x)=q(x),\,\|T\|_{{\mathrm{cb}}}\leq 1\}.$
    We now show that the expression on the right equals $t!$ times $\|T_{p}\|_{{\mathrm{cb}},*}$, which we recall can be written as
    (26) $\|T_{p}\|_{{\mathrm{cb}},*}=\sup\{\langle T_{p},T\rangle:\|T\|_{{\mathrm{cb}}}\leq 1\}.$
    By inclusion of the feasible region we have that $\|p\|_{{\mathrm{cb}},*}\leq t!\|T_{p}\|_{{\mathrm{cb}},*}$. For the other inequality, let $T\in\mathbb{R}^{n\times\dots\times n}$ be a $t$-tensor and consider $\Pi_{\mathcal{P}}T$ as in Eq. 10. By Proposition 3.7 we have $\|\Pi_{\mathcal{P}}T\|_{{\mathrm{cb}}}\leq\|T\|_{{\mathrm{cb}}}\leq 1$. Also note that the polynomial $\Pi_{\mathcal{P}}T(x)$ belongs to $V_{\mathcal{P}}$, because $(\Pi_{\mathcal{P}}T)_{\mathbf{i}}=0$ unless $\mathbf{i}$ contains exactly one index from each set in the partition $\mathcal{P}$. It remains to observe that $(\Pi_{\mathcal{P}}T)_{\mathbf{i}}=T_{\mathbf{i}}$ for all indices $\mathbf{i}\in[n]^{t}$ for which $(T_{p})_{\mathbf{i}}\neq 0$ and therefore
    $\langle T_{p},T\rangle=\langle T_{p},\Pi_{\mathcal{P}}T\rangle.$
    This shows that $\|p\|_{{\mathrm{cb}},*}\geq t!\|T_{p}\|_{{\mathrm{cb}},*}$. $\Box$
5.3. Alternative proof of the main result via semidefinite programming
First of all, we will state Theorem 5.5 (which corresponds, after some reformulation, to equation (20) of [GL19]), that gives an optimization problem equivalent to the dual of the SDP $\mathcal{E}(p,t)$. Before that, we introduce the following notation. Given $\mathbf{i}\in[n+1]^{2t}$, $\alpha(\mathbf{i})\in\{0,1\}^{n}$ is defined as
$(\alpha(\mathbf{i}))_{m}:=\left\{\begin{array}[]{ll}1&\mathrm{if}\ m\in[n]\ \mathrm{and}\ m\ \mathrm{occurs\ an\ odd\ number\ of\ times\ in\ }\mathbf{i},\\ 0&\mathrm{otherwise}.\end{array}\right.$
Theorem 5.5 ([GL19]).
Let $p:\{-1,1\}^{n}\to\mathbb{R}$ and $t\in\mathbb{N}$. Then,
(27) $\mathcal{E}(p,t)=\sup$ $\ (\langle p,r\rangle-w)/\|r\|_{1}$
$\mathrm{s.t.}$ $\ r:\{-1,1\}^{n}\to\mathbb{R},\ d\in\mathbb{R}$
$\ A_{s}\colon[n+1]\to B_{M(d)}\ \mathrm{for\ all}\ s\in[2t]$
$\ u,v\in\mathbb{R}^{d},\ w=\|u\|^{2}=\|v\|^{2}$
$\ c_{\alpha(\mathbf{i})}=\langle u,A_{1}(i_{1})\dots A_{2t}(i_{2t})v\rangle\ \mathrm{for\ all}\ \mathbf{i}\in[n+1]^{2t},$
where $c_{\alpha}$ are the coefficients of $r$ [^4] Following [GL19] one obtains Theorem 5.5, but with the $A_{s}$ being unitary-valued maps. Every contraction-valued map can be turned into an equivalent unitary-valued map by block-encoding contractions into the top-left corner of unitaries.
Second, we show that in the case of $p$ belonging to $W_{\mathcal{P}}$, we can restrict $r$ to belong to $W_{\mathcal{P}}$. As before, we also show that a single contraction-valued map $A$ suffices. [^5] In fact, all of the contraction-valued maps $A_{s}$ can be taken to be the same regardless $p$ belongs to $W_{\mathcal{P}}$ or not, similarly to what is done in Proposition 3.2.
Lemma 5.6.
Let $\mathcal{P}$ be a partition of $[n]$ in $2t$ subsets and let $p\in W_{\mathcal{P}}$. Then,
(28) $\mathcal{E}(p,t)=\sup$ $\ (\langle p,r\rangle-w)/\|r\|_{1}$
$\mathrm{s.t.}$ $\ r:\{-1,1\}^{n}\to\mathbb{R},\,r\in W_{\mathcal{P}},\,d\in\mathbb{R}$
$\ A\colon[n+1]\to B_{M(d)},\,\mathrm{for}\ s\in[2t]$
$\ u,v\in\mathbb{R}^{d},\,w=\|u\|^{2}=\|v\|^{2}$
$\ c_{\alpha(\mathbf{i})}=\langle u,A(\mathbf{i})v\rangle\ \mathrm{for\ all}\ \mathbf{i}\in[n+1]^{2t}\ ,$
where $c_{\alpha}$ are the coefficients of $r$.
  • Proof:
    Let *$\mathcal{E}^{*}(p,t)$ be the expression in the right-hand side of Eq. 28. By inclusion of the feasible region, $\mathcal{E}(p,t)\geq\mathcal{E}^{*}(p,t)$. To prove the other inequality, consider a feasible instance $(u,v,w,A_{s},r)$ for the SDP (27). Consider the contraction-valued map $\tilde{A}(i):=\sum_{s\in[2t]}e_{s}e_{s+1}^{\mathsf{T}}\otimes\hat{A}_{s}(i)$ for $i\in[n]$, where $\hat{A}_{s}(i)$ is a $d2^{2t}\times d2^{2t}$ matrix defined by
    $\hat{A}_{s}(i)=\bigoplus_{z\in\{-1,1\}^{2t}}(A_{s}\cdot z)(i).$
    We also define $\tilde{A}(n+1):=0$ and the vectors $\tilde{u}=e_{1}\otimes\hat{u}$ and $\tilde{v}=e_{2t+1}\otimes\hat{v}$, where $\hat{u}$ is the $2^{t}d$-dimensional vector defined as the (normalized) direct sum of $2^{t}$ copies of $u$, i.e.,
    $\hat{u}=\frac{1}{\sqrt{2^{2t}}}\bigoplus_{z\in\{-1,1\}^{2t}}u,$
    and the same for $\hat{v}$, but with an appropriate sign in each of the copies
    $\hat{v}=\frac{1}{\sqrt{2^{2t}}}\bigoplus_{z\in\{-1,1\}^{2t}}v\prod_{I\in\mathcal{P}}z_{I}.$
    This way, $(\tilde{u},\tilde{v},w,\tilde{A},\Pi_{\mathcal{P}}r)$ is a feasible instance of *$\mathcal{E}^{*}(p,t).$ Indeed, if $\mathbf{i}$ takes the value $n+1$ at least once or has any repeated indices, then $\alpha(\mathbf{i})_{1}+\dots+\alpha(\mathbf{i})_{n}<2t$, so $(\Pi_{\mathcal{P}}r)_{\alpha(\mathbf{i})}=0$ because $\Pi_{\mathcal{P}}r\in W_{\mathcal{P}}$, and $\langle\tilde{u},\tilde{A}(\mathbf{i})\tilde{v}\rangle=0$ by construction. If $\mathbf{i}\in[n]^{2t}$ and has no repeated indices, we get that
    $\langle\tilde{u},\tilde{A}(\mathbf{i})\tilde{v}\rangle=\frac{1}{2^{2t}}\sum_{z\in\{-1,1\}^{2t}}\langle u,A_{1}(i_{1})\dots A_{2t}(i_{2t})v\rangle\prod_{I\in\mathcal{P}}z_{I}^{1+\sum_{j\in I}\alpha(\mathbf{i})_{j}}.$
    Now, reasoning as in Proposition 3.7, we get that
    $\langle\tilde{u},\tilde{A}(\mathbf{i})\tilde{v}\rangle=\left\{\begin{array}[]{ll}\langle u,A_{1}(i_{1})\dots A_{2t}(i_{2t})v\rangle&\mathrm{if}\ \mathbf{i}\ \text{takes one value in each }I\in\mathcal{P},\\ 0&\mathrm{otherwise},\end{array}\right.$
    so putting everything together we get that
    $\tilde{c}_{\alpha(\mathbf{i})}=\langle\tilde{u},\tilde{A}(\mathbf{i})\tilde{v}\rangle,$
    where $\tilde{c}_{\alpha}$ are the coefficients of $\Pi_{\mathcal{P}}r$. Since $\tilde{A}$ is contraction-valued and both $\|\tilde{u}\|^{2}=\frac{1}{2^{2t}}\sum_{z\in\{-1,1\}^{2t}}\|u\|^{2}=\|u\|^{2}=w$ and $\|\tilde{v}\|^{2}=w$, we conclude that $(\tilde{u},\tilde{v},w,\tilde{A},\Pi_{\mathcal{P}}r)$ is a feasible instance of Eq. 28. Finally, the value of $(\tilde{u},\tilde{v},w,\tilde{A},\Pi_{\mathcal{P}}r)$ is at least as large as the one of $(u,v,w,A,r)$:
    $\frac{\langle p,\Pi_{\mathcal{P}}r\rangle-w}{\|\Pi_{\mathcal{P}}r\|_{1}}=\frac{\langle p,r\rangle-w}{\|\Pi_{\mathcal{P}}r\|_{1}}\geq\frac{\langle p,r\rangle-w}{\|r\|_{1}},$
    where in the equality we have used that that $p$ belongs to $W_{\mathcal{P}}$ and in the inequality we have used Lemma 3.8. $\Box$
Now, we are ready to prove Theorem 5.1, again.
  • Proof of Theorem 5.1:
    First note that given a feasible instance $(u,v,w,A,r)$ for Eq. 28 we clearly have that $r_{=2t}\in V_{\mathcal{P}}$. We show that also $w\geq\|r_{=2t}\|_{{\mathrm{cb}},*}$. By Propositions 3.3 and 5.4, this requires us to show that $t!$ times the unique symmetric $2t$-tensor $T_{r_{=2t}}$ associated to $r_{=2t}$ belongs to $wK(n,t)$. To do so, we show that $t!T_{r_{=2t}}=(\langle u,A(\mathbf{i})v\rangle)_{\mathbf{i}\in[n]^{2t}}$. Let $\mathbf{i}\in[n]^{2t}$. If $\mathbf{i}$ has repeated indices then $(T_{r_{=2t}})_{\mathbf{i}}=0$ because $r_{=2t}$ is multilinear, and also $\langle u,A(\mathbf{i})v\rangle=0$, because $r$ is a feasible instance of Eq. 28. If $\mathbf{i}$ does not have repeated indices, then $t!(T_{r_{=2t}})_{\mathbf{i}}=c_{\alpha(\mathbf{i})}$, and also $\langle u,A(\mathbf{i})v\rangle=c_{\alpha(\mathbf{i})}$ because $r$ is a feasible solution for Eq. (28). On the other hand, given $r\in W_{\mathcal{P}}$ there is an instance $(u,v,w,A,r)$ with $w=\|r_{=2t}\|_{{\mathrm{cb}},*}$. Indeed, by Proposition 5.4 there is a map $A\colon[n]\to B_{M(d)}$ and vectors $u,v$ whose norm squared is $\|r_{=2t}\|_{{\mathrm{cb}},*}$ such that $c_{\alpha}(r_{=2t})=\langle u,A(\mathbf{i})v\rangle$ for every $\mathbf{i}\in\mathcal{I}_{\alpha}$ and every $\alpha\in\mathbb{Z}_{\geq 0}^{n}$. Note that in order to have a feasible instance for Eq. 28 we need to satisfy its last condition, and with these contractions we can only satisfy it for $\alpha$ such that $\alpha_{1}+\dots+\alpha_{n}=2t$. To satisfy it for every $\alpha$ with $\alpha_{1}+\dots+\alpha_{n}\leq 2t$, we just have to change the contractions and the vectors by $\hat{A}(i):=\sum_{s\in[2t]}e_{s}e_{s+1}^{\mathsf{T}}\otimes A_{s}(i)$ and $\hat{u}:=e_{1}\otimes u$ and $\hat{v}:=e_{t+1}\otimes v$ and define the extra contraction as $\hat{A}(n+1)=0$. This, way $(\hat{u},\hat{v},\|r_{=2t}\|_{{\mathrm{cb}},*},\hat{A},r)$ is a feasible instance. To sum up, so far we have proved that
    $\mathcal{E}(p,t)=\sup$ $\ (\langle p,r\rangle-\|r_{=2t}\|_{{\mathrm{cb}},*})/\|r\|_{1}$
    $\mathrm{s.t.}$ $\ r:\{-1,1\}^{n}\to\mathbb{R},\ r\in W_{\mathcal{P}}.$
    We finally reformulate the above in terms of $r_{=2t}$ using the following two observations. Since $p\in V_{\mathcal{P}}$ we have $\langle p,r\rangle=\langle p,r_{=2t}\rangle$. Moreover, by Proposition 5.2, we have $\|r_{=2t}\|_{\infty,*}=\inf\{\|\tilde{r}\|_{1}\,\mid\,\tilde{r}:\{-1,1\}^{n}\to\mathbb{R},\,\tilde{r}\in W_{\mathcal{P}},\,\tilde{r}_{=2t}=r_{=2t}\}$. Hence,
    $\mathcal{E}(p,t)=\sup$ $\ (\langle p,r_{=2t}\rangle-\|r_{=2t}\|_{{\mathrm{cb}},*})/\|r_{=2t}\|_{\infty,*}$
    $\mathrm{s.t.}$ $\ r_{=2t}\in V_{\mathcal{P}},$
    which concludes the proof. $\Box$
6. No additive converse for the polynomial method
In this section we will show that there is no analogue of Corollary 1.2 for polynomials of degree 4. In other words, we will prove Theorem 1.4.
See 1.4
  • Proof:
    For any partition $\mathcal{P}$ of $\{0\}\cup[3n]$ in $2t$ subsets, Theorem 5.1 shows that
    $\sup_{p\in V_{\mathcal{P}},\|p\|_{\infty}\leq 1}\mathcal{E}(p,t)$ $=\sup_{p\in V_{\mathcal{P}},\|p\|_{\infty}\leq 1}\sup_{r\in V_{\mathcal{P}},\|r\|_{\infty,*}\leq 1}\langle p,r\rangle-\|r\|_{{\mathrm{cb}},*}$
    $=\sup_{r\in V_{\mathcal{P}},\|r\|_{\infty,*}\leq 1}\|r\|_{\infty,*}-\|r\|_{{\mathrm{cb}},*}$
    $=\sup_{r\in V_{\mathcal{P}},\|r\|_{\infty,*}=1}1-\|r\|_{{\mathrm{cb}},*}.$
    Consider now the case $t=2$ and the partition $\mathcal{P}_{n}=\{\{0\},\{1,\dots,n\},\{n+1,\dots,2n\},\{2n+1,\dots,3n\}\}$ of $\{0\}\cup[3n]$. In Section 4.4 a sequence of forms $p_{n}\in V_{\mathcal{P}_{n}}$ was constructed with the property that
    (29) $\frac{\|p_{n}\|_{{\mathrm{cb}}}}{\|p_{n}\|_{\infty}}\to\infty.$
    Hence, by a duality argument we get that there is a sequence $r_{n}\in V_{\mathcal{P}_{n}}$ such that $\|r_{n}\|_{{\mathrm{cb}},*}/\|r_{n}\|_{\infty,*}\to 0$. Indeed, suppose towards a contradiction that there is a $K>0$ such that for every $n\in\mathbb{N}$ and every $r\in V_{\mathcal{P}_{n}}$ we have that $\|r\|_{{\mathrm{cb}},*}\geq K\|r\|_{\infty,*}$. Then,
    $\|p\|_{{\mathrm{cb}}}=\sup_{\|r\|_{{\mathrm{cb}},*}\leq 1}\langle r,p\rangle\leq\frac{1}{K}\sup_{\|r\|_{\infty,*}\leq 1}\langle r,p\rangle=\frac{1}{K}\|p\|_{\infty},$
    which contradicts Eq. 29. The sequence $r_{n}$ shows that
    $\sup_{p\in V_{\mathcal{P}_{n}},\|p\|_{\infty}\leq 1,n\in\mathbb{N}}\mathcal{E}(p,2)=1,$
    which implies the stated result. $\Box$
7. Characterizing $K_{G}$ with $1$-query quantum algorithms
In this section we show that the best additive error up to which 1-query quantum algorithms can compute bounded bilinear forms is exactly $1-1/K_{G}$, where $K_{G}$ is the real Grothendieck constant.
7.1. Warm up: the case $n=2$ via the CHSH game.
The celebrated CHSH game defines a bilinear form that saturates Grothendieck’s theorem for $2\times 2$ matrices [CHSH69]. In other words,
$K_{G}(2):=\sup\left\{\frac{\|A\|_{\mathrm{cb}}}{\|A\|_{\infty}}:A\in\mathbb{R}^{2\times 2}\right\}$
is the $\sqrt{2}$ attained by the CHSH bilinear form [FR94]. In the language of XOR games, the CHSH game is known to attain Tsirelson’s bound for games with two questions. That is, $\sqrt{2}$ is the biggest quantum bias-classical bias ratio for XOR games of two questions. Indeed, phrased in terms of norms, Tsirelson showed that every XOR game $G$ determines a matrix $A$ such that the quantum bias of $G$ equals $\|A\|_{{\mathrm{cb}}}$ and the classical bias equals $\|A\|_{\infty}$ [Tsi80].
Here, we use the CHSH bilinear form to prove the analogue of Theorem 1.5 restricted to $n=2$ without using our main result Theorem 5.1.
Theorem 7.1.
The worst-case minimum error for one-query quantum algorithms satisfies
$\sup_{p}\mathcal{E}(p,1)=1-\frac{1}{\sqrt{2}},$
where the supremum is taken over the set of bounded bilinear forms in $4$ variables.
  • Proof:
    The upper bound was proven in [${}^{+}$16]. For the lower bound consider the bilinear form associated to the CHSH game $f:\{-1,1\}^{4}\to\{-1,1\}$:
    $f(x)=\frac{1}{2}\big{(}x_{1}(x_{3}+x_{4})+x_{2}(x_{3}-x_{4})\big{)}=\frac{1}{2}(x_{1}\ x_{2})\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\begin{pmatrix}x_{3}\\ x_{4}\end{pmatrix}.$
    One can verify that $f(x)\in\{-1,1\}$ for all $x\in\{-1,1\}^{4}$ and therefore $\|f\|_{1}=\|f\|_{2}^{2}=1$. We now emulate the construction from Lemma 4.2. Write the coefficients of $f$ as $c_{S}$ for $S\in{[4]\choose 2}$. Then, for each $i\in[4]$ define the unit vector $w_{i}\in\mathbb{R}^{4}$ by
    $w_{i}=\sqrt{2}\sum_{j\in[4]\setminus\{i\}}c_{\{i,j\}}e_{j}.$
    Now define the matrices $A:[n]\to B_{M(6)}$ by
    $A(i)=\begin{pmatrix}0&0&0\\ w_{i}&0&0\\ 0&e_{i}^{\mathsf{T}}&0\end{pmatrix}$
    It is easily verified that $A(i)^{2}=0$ and that the $(6,1)$-coordinate of $A(i)A(j)$ equals $\sqrt{2}c_{\{i,j\}}$ if $i\neq j$, from which it also follows that these matrices commute. Setting $A(5)=0$, we get that
    $\langle e_{6},A(i)A(j)e_{1}\rangle=\begin{cases}\sqrt{2}c_{\{i,j\}}&\text{if $\{i,j\}\in{[4]\choose 2}$}\\ 0&\text{otherwise.}\end{cases}$
    Defining $r=\sqrt{2}f$, it follows that $A,u,v$ and $r$ define a feasible solution for the optimization problem of Lemma 5.6, which attains the value
    $\mathcal{E}(f,1)\geq(\langle f,\sqrt{2}f\rangle-1)/\sqrt{2}=1-\frac{1}{\sqrt{2}}.$
    $\Box$
7.2. The general case
We recall that it was shown in [${}^{+}$16] that for every bilinear form there exists a 1-query quantum algorithm that makes additive error at most $1-1/K_{G}$. It thus remains to show the lower bound. See 1.5
  • Proof:
    Theorem 5.1 shows the following:
    (30) $\sup_{p\in\mathcal{BB}}\mathcal{E}(p,1)$ $=\sup_{\|p\|_{\infty}\leq 1}\sup_{\|r\|_{\infty,*}\leq 1}\langle p,r\rangle-\|r\|_{{\mathrm{cb}},*}$
    (31) $=\sup_{\|r\|_{\infty,*}\leq 1}\|r\|_{\infty,*}-\|r\|_{{\mathrm{cb}},*}$
    $=\sup_{\|r\|_{\infty,*}=1}1-\|r\|_{{\mathrm{cb}},*}.$
    It thus remains to show that for bilinear forms $\|r\|_{\infty,*}\leq K_{G}\|r\|_{{\mathrm{cb}},*}$. We do so starting from Grothendieck’s theorem for matrices. It states that for $A\in\mathbb{R}^{n\times n}$ we have $\|A\|_{{\mathrm{cb}}}\leq K_{G}\|A\|_{\infty}$. Each bilinear form $q:\{-1,1\}^{n}\times\{-1,1\}^{n}\to\mathbb{R}$ uniquely corresponds to a matrix $A\in\mathbb{R}^{n\times n}$ such that $q(x,y)=x^{\mathsf{T}}Ay$. Moreover, for such $q$ and $A$ one has $\|q\|_{\infty}=\|A\|_{\infty}$ (immediate) and in Corollary 3.6 we showed $\|q\|_{{\mathrm{cb}}}=\|A\|_{{\mathrm{cb}}}$, so $\|q\|_{\mathrm{cb}}\leq K_{G}\|q\|_{\infty}$. A duality argument then concludes the proof:
    $\|r\|_{\infty,*}=\sup_{\|q\|_{\infty}\leq 1}\langle r,q\rangle\leq\sup_{\|q\|_{{\mathrm{cb}}}\leq K_{G}}\langle r,q\rangle=K_{G}\|r\|_{{\mathrm{cb}},*}.$
    $\Box$
Remark 7.2.
If in Theorem 1.5 we restrict the supremum to bilinear forms on $n+n$ variables, for a fixed $n$, then we obtain a characterization of $K_{G}(n)$ instead of $K_{G}$. Here, $K_{G}(n)=\sup\|A\|_{{\mathrm{cb}}}/\|A\|_{\infty}$, where the supremum is taken over all non-zero $n\times n$ real matrices.
8. An open question
Let $\mathcal{P}$ be a partition of $[n]$ in $2t$ subsets and let $p\in V_{\mathcal{P}}$ with $\|p\|_{\infty}\leq 1$. From the characterization of quantum $t$-query algorithms of [ABP19] we know that there is a quantum query algorithm $\mathcal{A}$ that outputs $p/\|p\|_{{\mathrm{cb}}}$ on expectation. In particular,
$|\mathbb{E}[\mathcal{A}(x)]-p(x)|=\left|\frac{p(x)}{\|p\|_{{\mathrm{cb}}}}-p(x)\right|\leq\|p\|_{\infty}\Big{(}1-\frac{1}{\|p\|_{{\mathrm{cb}}}}\Big{)}.$
As a consequence, one has that
(32) $\mathcal{E}(p,t)\leq\|p\|_{\infty}\Big{(}1-\frac{1}{\|p\|_{{\mathrm{cb}}}}\Big{)}.$
Our Theorem 5.1 implies that both sides of Equation (32) are equal when you take the supremum over all $p\in V_{\mathcal{P}}$. We wonder if that is true for every $p\in V_{\mathcal{P}}$.
Question 8.1.
Let $\mathcal{P}$ be a partition of $[n]$ in $2t$ subsets and let $p\in V_{\mathcal{P}}$. Is it true that
$\mathcal{E}(p,t)=\|p\|_{\infty}\Big{(}1-\frac{1}{\|p\|_{{\mathrm{cb}}}}\Big{)}?$
A positive answer to this question would strengthen our main technical contribution, Theorem 5.1. In addition, if we focus on the case $t=1$, it would imply that the method proposed in [${}^{+}$16] to give an algorithm that computes $p/\|p\|_{{\mathrm{cb}}}$ with $1$ query, provides the best $1$ query approximation for every $p$ (here the best means the one that minimizes $\mathcal{E}(p,1)$). For $t\geq 1$ it was showed in [ABP19] that $p/\|p\|_{\mathrm{cb}}$ is the output of a $t$ query algorithm, which would also be the best possible $t$ query approximation for $p$.
Finally, for the case $t=1$, it would imply a clean link between the biases of two player XOR games and quantum query algorithms. Indeed, given a matrix $A\in\mathbb{R}^{n\times n}$ it both defines a bounded bilinear form $p_{A}(x,y)=x^{\mathsf{T}}Ay$ and a two player XOR game $G_{A}$, where the referee asks the pair of questions $(i,j)$ with probability
$\pi(i,j)=\frac{|A_{i,j}|}{{\sum_{i,j\in[n]}|A_{i,j}|}}$
and the payoff is given by
$\mu(i,j,a,b)=\frac{1+ab\cdot\mathrm{sgn}[A_{i,j}]}{2}.$
Corollary 3.6 states that
$\|p_{A}\|_{\infty}=\|A\|_{\infty}\ \mathrm{and}\ \|p_{A}\|_{{\mathrm{cb}}}=\|A\|_{{\mathrm{cb}}},$
while Tsirelson’s work [Tsi80] implies that the classical and quantum biases of $G_{A}$ are
$\beta(G_{A})=\|A\|_{\infty}\ \mathrm{and}\ \beta^{*}(G_{A})=\|A\|_{{\mathrm{cb}}}.$
Thus, a positive answer to 8.1 would imply that
$\mathcal{E}(p_{A},1)=\beta(G_{A})\Big{(}1-\frac{1}{\beta^{*}(G_{A})}\Big{)}.$
Acknowledgments
We thank anonymous referees for their helpful feedback.